OpenStudy (anonymous):

what is the derivative of (2x+7)/(6x+2)

OpenStudy (anonymous):

first multiply them all together 12x^2 + 4x + 42x + 14 then take derivative 24x + 4 + 42 -> 24x + 44

sam (.sam.):

No you cant multiply its in fraction

OpenStudy (anonymous):

use the quotient rule here

OpenStudy (anonymous):

never mind, i thought that multiplification, my bad

OpenStudy (anonymous):

\[\frac{d}{dx}(\frac{2x+7)}{6x+2})=\frac{(6x+2)2-(2x+7)6}{(6x+2)^2}\] you can simplfiy

OpenStudy (anonymous):

the 2 and 6 in the numerator are not exponents, btw

OpenStudy (anonymous):

quotient rule

OpenStudy (anonymous):

(deniminator * derivative of numerator - numerator* derivative of denominator)/(denominator*benominator)

OpenStudy (anonymous):

denominator lol